# Kandinsky Conformal Prediction:

### v1

@@ A.1 Proofs from Section 3
Let $\alpha,\delta,{\cal W}$ be specified as in Theorem 3.1. Assume that $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, there exists an absolute constant $C$ such that the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\mathbb{E}_{D,E}\left[\sup_{w\in{\cal W}}\mathbb{E}_{(X_{n+1},Y_{n+1})\sim{\cal D},\varepsilon_{n+1}\sim{\cal D}_{\text{rn}}}\left[\frac{w(X_{n+1},Y_{n+1})}{\|w\|_{\infty}}\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1})\right\}-(1-\alpha)\right)\right]\right]\leq C\sqrt{\frac{d}{n}}+\frac{d}{n},$
@@
$\sup_{w\in{\cal W}}\mathbb{E}_{(X_{n+1},Y_{n+1})\sim{\cal D},\varepsilon_{n+1}\sim{\cal D}_{\text{rn}}}\left[\frac{w(X_{n+1},Y_{n+1})}{\|w\|_{\infty}}\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1})\right\}-(1-\alpha)\right)\right]\leq$
$\sup_{w\in{\cal W}}\left|\mathbb{E}_{(X_{n+1},Y_{n+1})\sim{\cal D},\varepsilon_{n+1}\sim{\cal D}_{\text{rn}}}\left[\frac{w(X_{n+1},Y_{n+1})}{\|w\|_{\infty}}\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1})\right\}-(1-\alpha)\right)\right]\right|\leq$
$C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}.$
Let the random variable $M_{D,E}=\sup_{w\in{\cal W}}\mathbb{E}_{(X_{n+1},Y_{n+1})\sim{\cal D},\varepsilon_{n+1}\sim{\cal D}_{\text{rn}}}\left[\frac{w(X_{n+1},Y_{n+1})}{\|w\|_{\infty}}\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1})\right\}-(1-\alpha)\right)\right]$. We have
$\mathbb{P}_{D,E}\left[M_{D,E}>C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right]\leq\delta.$
By inverting the bound, for $t>C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}$,
$\mathbb{P}_{D,E}\left[M_{D,E}>t\right]\leq 4\mathop{\rm exp}\left[-\frac{n\big{(}t-C^{\prime}\sqrt{\frac{d}{n}}-\frac{d}{n}\big{)}^{2}}{2\max\{\alpha,1-\alpha\}^{2}}\right].$
Since $M_{D,E}\geq 0$,
@@
$=\int_{t=0}^{C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}}\mathbb{P}_{D,E}[M_{D,E}>t]\,\mathrm{d}t+\int_{t=C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}}^{+\infty}\mathbb{P}_{D\sim{\cal D}^{n}}\left[M_{D}>t\right]\,\mathrm{d}t$
$\leq C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\int_{t=C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}}^{+\infty}\mathbb{P}_{D\sim{\cal D}^{n}}\left[M_{D}>t\right]\,\mathrm{d}t$
$\leq C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\int_{t=C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}}^{+\infty}4\mathop{\rm exp}\left[-\frac{n\big{(}t-C^{\prime}\sqrt{\frac{d}{n}}-\frac{d}{n}\big{)}^{2}}{2\max\{\alpha,1-\alpha\}^{2}}\right]\,\mathrm{d}t$
$=C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{8\pi}{n}}$
$\leq\left(C^{\prime}+\sqrt{8\pi}\right)\sqrt{\frac{d}{n}}+\frac{d}{n}.$
@@
$\left|\mathbb{P}\left[Y_{n+1}\in{\cal C}(X_{n+1})\mid Z_{n+1}\in G\right]-(1-\alpha)\right|\leq$
$\frac{1}{\mathbb{P}[Z\in G]}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right).$
@@
According to Theorem 3.1, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness in $\{(X_{i},Y_{i})\}_{i\in[n]}$ and $\{\varepsilon_{i}\}_{i\in[n]}$, for every $w\in{\cal W}$,
$\left|\mathbb{E}_{(X_{n+1},Y_{n+1})\varepsilon_{n+1}}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{{\widetilde{S}}(X_{n+1},Y_{n+1},\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\right)\right]\right|=$
$\left|\mathbb{E}_{(X_{n+1},Y_{n+1})\varepsilon_{n+1}}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1})\right\}-(1-\alpha)\right)\right]\right|\leq$
${}\|w\|_{\infty}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right)=$
$\|w\|_{\infty}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right).$
For any $G\in{\cal G}$, by taking $w(x,y)=\frac{1}{\mathbb{P}[Z\in G]}\mathbb{P}[Z\in G\mid\phi(X,Y)=\phi(x,y)]$, we have $\|w\|_{\infty}\leq\frac{1}{\mathbb{P}[Z\in G]}$. Since ${\widetilde{S}}(X_{n+1},Y_{n+1},\varepsilon_{n+1})$, $\widehat{q}(X_{n+1},Y_{n+1})$ are measurable w.r.t. $\phi(X_{n+1},Y_{n+1}),\varepsilon_{n+1}$, by Bayes formula,
$\mathbb{E}_{(X_{n+1},Y_{n+1}),\varepsilon_{n+1}}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{{\widetilde{S}}(X_{n+1},Y_{n+1},\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\right)\right]=$
$\mathbb{E}_{(X_{n+1},Y_{n+1}),\varepsilon_{n+1}}\left[\frac{\mathbb{P}[Z_{n+1}\in G\mid\phi(X_{n+1},Y_{n+1})]}{\mathbb{P}[Z_{n+1}\in G]}\left(\mathbbm{1}\left\{{\widetilde{S}}(X_{n+1},Y_{n+1},\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\right)\right]=$
$\mathbb{E}_{\phi(X_{n+1},Y_{n+1}),\varepsilon_{n+1}}\Big{[}\frac{\mathbb{P}[Z_{n+1}\in G\mid\phi(X_{n+1},Y_{n+1}),\varepsilon_{n+1}]}{\mathbb{P}[Z_{n+1}\in G\mid\varepsilon_{n+1}]}\cdot$
$\hskip 130.08731pt\left(\mathbbm{1}\left\{{\widetilde{S}}(X_{n+1},Y_{n+1},\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\right)\Big{]}=$
$\mathbb{E}_{(X_{n+1},Y_{n+1}),\varepsilon_{n+1}}\left[\mathbbm{1}\left\{{\widetilde{S}}(X_{n+1},Y_{n+1},\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\Bigm{|}Z_{n+1}\in G\right]=$
$\mathbb{P}_{(X_{n+1},Y_{n+1}),\varepsilon_{n+1}}\left[Y_{n+1}\in{\cal C}(X_{n+1})\mid Z_{n+1}\in G\right]-(1-\alpha).$
@@
$\left|\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1})\mid(X_{n+1},Y_{n+1})\in G]-(1-\alpha)\right|\leq$
${}\frac{1}{\mathbb{P}[(X_{n+1},Y_{n+1})\in G]}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right).$
@@
Let $\alpha$ and $\delta$ be parameters in $(0,1)$, and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$|\mathbb{P}_{(X_{n+1},Y_{n+1})\sim{\cal D}_{\text{T}},\varepsilon_{n+1}}[Y_{n+1}\in{\cal C}(X_{n+1})]-(1-\alpha)|\leq$
$B\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right),$
where $(X_{n+1},Y_{n+1})$ are drawn independently from a distribution ${\cal D}_{\text{T}}$, for every distribution ${\cal D}_{\text{T}}$ such that $\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}\in{\cal W}$ and $\left|\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}(x,y)\right|\leq B$ for any $x\in{\cal X}$ and $y\in{\cal Y}$.
@@
This corollary follows from Theorem 3.1 for any weight function $w\in{\cal W}$ of the form $w(x,y)=\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}(x,y)$ with $\|w\|_{\infty}\leq B$. ∎

### v2

@@ A.1 Proofs from Section 3
Let $\alpha,\delta,{\cal W},w_{\beta}$ be specified as in Theorem 3.1. Assume that $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, there exists an absolute constant $C$ such that the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\mathbb{E}_{D,E}\left[\sup_{w_{\beta}\in{\cal W}}\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[\frac{w_{\beta}(X,Y)}{\|\beta\|_{1}}\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right]\leq C\sqrt{\frac{d}{n}}+\frac{d}{n},$
@@
$\sup_{w_{\beta}\in{\cal W}}\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[\frac{w_{\beta}(X,Y)}{\|\beta\|_{1}}\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\leq$
$\sup_{w_{\beta}\in{\cal W}}\left|\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[\frac{w_{\beta}(X,Y)}{\|\beta\|_{1}}\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq$
$C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}\leq$
$C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}.$
Let the random variable $M_{D,E}=\sup_{w_{\beta}\in{\cal W}}\mathbb{E}_{(X,Y)\sim{\cal D},\varepsilon\sim{\cal D}_{\text{rn}}}\left[\frac{w_{\beta}(X,Y)}{\|\beta\|_{1}}\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]$. We have
$\mathbb{P}_{D,E}\left[M_{D,E}>C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right]\leq\delta.$
By inverting the bound, for $t>C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}$,
$\mathbb{P}_{D,E}\left[M_{D,E}>t\right]\leq 4\mathop{\rm exp}\left[-\frac{n\big{(}t-C^{\prime}\sqrt{\frac{d}{n}}-\frac{d}{n}-\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}\big{)}^{2}}{2\max\{\alpha,1-\alpha\}^{2}}\right].$
Let $I=C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}$. Since $M_{D,E}\geq 0$,
@@
$=\int_{t=0}^{I}\mathbb{P}_{D,E}[M_{D,E}>t]\,\mathrm{d}t+\int_{t=I}^{+\infty}\mathbb{P}_{D,E}[M_{D,E}>t]\,\mathrm{d}t$
$\leq C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}+\int_{t=I}^{+\infty}\mathbb{P}_{D,E}[M_{D,E}>t]\,\mathrm{d}t$
$\leq C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}+\int_{t=I}^{+\infty}4\mathop{\rm exp}\left[-\frac{n\big{(}t-I\big{)}^{2}}{2\max\{\alpha,1-\alpha\}^{2}}\right]\,\mathrm{d}t$
$=C^{\prime}\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln d}{n}}+\max\{\alpha,1-\alpha\}\sqrt{\frac{8\pi}{n}}$
$\leq\left(C^{\prime}+\sqrt{2}+\sqrt{8\pi}\right)\sqrt{\frac{d}{n}}+\frac{d}{n}.$
@@
$\left|\mathbb{P}\left[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid Z_{n+1}\in G,\{X_{i},Y_{i},\varepsilon_{i}\}_{i\in[n]}\right]-(1-\alpha)\right|\leq$
$\frac{1}{\mathbb{P}[Z_{n+1}\in G]}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4|{\cal G}|/\delta)}{n}}\right).$
@@
According to Theorem 3.1, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness in $\{(X_{i},Y_{i})\}_{i\in[n]}$ and $\{\varepsilon_{i}\}_{i\in[n]}$, for every $w_{\beta}\in{\cal W}$,
$\left|\mathbb{E}_{X,Y,\varepsilon}\left[w_{\beta}(X,Y)\left(\mathbbm{1}\left\{{\widetilde{S}}(X,Y,\varepsilon)\leq\widehat{q}(X,Y)\right\}-(1-\alpha)\right)\right]\right|=$
$\left|\mathbb{E}_{X,Y,\varepsilon}\left[w_{\beta}(X,Y)\left(\mathbbm{1}\left\{Y\in{\cal C}(X;\varepsilon)\right\}-(1-\alpha)\right)\right]\right|\leq$
${}\|\beta\|_{1}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}\right)=$
$\|\beta\|_{1}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4|{\cal G}|/\delta)}{n}}\right).$
For any $G\in{\cal G}$, by taking $w_{\beta}(x,y)=\frac{1}{\mathbb{P}[Z\in G]}\mathbb{P}[Z\in G\mid\phi(X,Y)=\phi(x,y)]$, we have $\|\beta\|_{1}=\frac{1}{\mathbb{P}[Z_{n+1}\in G]}$. Since ${\widetilde{S}}(X,Y,\varepsilon)$, $\widehat{q}(X,Y)$ are measurable w.r.t. $\phi(X,Y),\varepsilon$, by Bayes formula,
$\mathbb{E}_{X,Y,\varepsilon}\left[w_{\beta}(X,Y)\left(\mathbbm{1}\left\{{\widetilde{S}}(X,Y,\varepsilon)\leq\widehat{q}(X,Y)\right\}-(1-\alpha)\right)\right]=$
$\mathbb{E}_{X,Y,\varepsilon}\left[\frac{\mathbb{P}[Z\in G\mid\phi(X,Y)]}{\mathbb{P}[Z\in G]}\left(\mathbbm{1}\left\{{\widetilde{S}}(X,Y,\varepsilon)\leq\widehat{q}(X,Y)\right\}-(1-\alpha)\right)\right]=$
$\mathbb{E}_{\phi(X,Y),\varepsilon}\Big{[}\frac{\mathbb{P}[Z\in G\mid\phi(X,Y),\varepsilon]}{\mathbb{P}[Z\in G\mid\varepsilon]}\cdot\left(\mathbbm{1}\left\{{\widetilde{S}}(X,Y,\varepsilon)\leq\widehat{q}(X,Y)\right\}-(1-\alpha)\right)\Big{]}=$
$\mathbb{E}_{\phi(X,Y),\varepsilon}\left[\mathbbm{1}\left\{{\widetilde{S}}(X,Y,\varepsilon)\leq\widehat{q}(X,Y)\right\}-(1-\alpha)\Bigm{|}Z\in G\right]=$
$\mathbb{E}\left[\mathbbm{1}\left\{{\widetilde{S}}(X_{n+1},Y_{n+1},\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},Y_{n+1})\right\}-(1-\alpha)\Bigm{|}Z_{n+1}\in G,\{X_{i},Y_{i},\varepsilon_{i}\}_{i\in[n]}\right]=$
$\mathbb{P}\left[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid Z_{n+1}\in G,\{X_{i},Y_{i},\varepsilon_{i}\}_{i\in[n]}\right]-(1-\alpha).$
@@
$\left|\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid(X_{n+1},Y_{n+1})\in G,\{X_{i},Y_{i},\varepsilon_{i}\}_{i\in[n]}]-(1-\alpha)\right|\leq$
${}\frac{1}{\mathbb{P}[(X_{n+1},Y_{n+1})\in G]}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4|{\cal G}|/\delta)}{n}}\right).$
@@
Let $\alpha$ and $\delta$ be parameters in (0,1)(0,1)( 0 , 1 ), and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, there exists an absolute constant $C$ such that, for every distribution ${\cal D}_{\text{T}}$ such that $\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}\in{\cal W}$ and $\left|\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}(x,y)\right|\leq B$ for any $x\in{\cal X}$ and $y\in{\cal Y}$, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$|\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid\{X_{i},Y_{i},\varepsilon_{i}\}_{i\in[n]}]-(1-\alpha)|\leq B\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right),$
where $(X_{n+1},Y_{n+1})$ are drawn independently from the distribution ${\cal D}_{\text{T}}$.
@@
This corollary follows from Lemma A.1 for any weight function $w\in{\cal W}$ of the form $w(x,y)=\frac{d\mathbb{P}_{{\cal D}_{T}}}{d\mathbb{P}_{{\cal D}}}(x,y)$ with $\|w\|_{\infty}\leq B$. ∎